OpenStudy (anonymous):

(1/4)^(2x+1) = 16 ^ (3x+2)

jhonyy9 (jhonyy9):

1/4= 2 on exponent minus 2 16= 2 on exponent 4 so hence we get -2(2x+1)=4(3x+2) -4x-2=12x+8 -10=16x x= -10/16 x= -5/8